# Chaos and spatial prethermalization in driven-dissipative bosonic chains

### v2

@@ 1. Numerical approaches
$\hat{L}_{1}=\hat{a}^{\dagger}(\gamma^{\uparrow}-\mathcal{S}\hat{a}\hat{a}^{\dagger})/\sqrt{\gamma^{\uparrow}},\qquad\hat{L}_{\phi}=\sqrt{\gamma^{\phi}}\hat{a}\hat{a}^{\dagger},\qquad\hat{L}_{\downarrow}=\sqrt{\gamma^{\downarrow}}\hat{a}.$ (39)
@@
while for the first model one requires $\gamma^{\rm s}\langle\hat{a}\hat{a}^{\dagger}\rangle\ll\gamma^{\uparrow}$. As discussed in the main text, the choice of $\hat{H}_{\rm imp}$ and $\gamma^{\phi}$ does not affect the steady-state properties, including the steady-state Wigner function $W(\alpha,\alpha^{*})$. In the steady state, they have an effect only on time-correlation functions. In particular, by imposing $\gamma^{\phi}=3\mathcal{S}/2$ one finds oscillating time-correlation functions with a usually very small decay rate (that is $23\mathcal{S}/$ 2 which is by assumption much smaller than $2\gamma^{\uparrow}/\langle\hat{a}\hat{a}^{\dagger}\rangle$). If instead $\gamma^{\phi}\gg\mathcal{S}$ one obtains the generalized version of the Scully-Lamb model exhibiting rapidly decaying time-correlation functions in the lasing phase. A similar phenomenology occurs in the 2-photon decay impurity model, where the phase coherence is suppressed if $\gamma^{\phi}\gg\gamma^{\rm s}$.
While the two impurity models above exhibits very similar steady-state properties, a finite (inverse) effective temperature $\beta^{\rm eff}$ can only be simply associated with the 2-photon decay model. Indeed, the set of Lindblad operators in Eq. (38) can be easily interpreted as coupling to a set of thermal reservoirs. $\hat{L}_{\rm s}$ corresponds to a zero-temperature 2-photon reservoir. $\hat{L}_{\uparrow}$ and $\hat{L}_{\downarrow}$ correspond to finite-temperature 1-photon reservoir at a temperature set by the detailed balance
$\frac{\gamma^{\uparrow}}{\gamma^{\downarrow}}=:{\rm e}^{-\beta^{\rm eff}(\omega_{0}-\mu)}=:{\rm e}^{\beta^{\rm eff}\mu^{\rm eff}},\quad\textit{i.e.}\quad\beta^{\rm eff}\mu^{\rm eff}:=\log\left({\gamma_{\uparrow}}\big{/}{\gamma_{\downarrow}}\right),$ (41)
@@
Figure 11: Fit of the $\ell$-th site Wigner function of the driven-dissipative bosonic chain described by Eqs. (2) and (1) for a chain with length $L=400$. In panel $(a)$, we show the $L^{2}$ norm in Eq. (42) between the $L$-th site Wigner function and the Scully-Lamb fitted Wigner function for one site each ten along the chain (green solid line). Additionally, we present the $L^{2}$ norms associated with the 2-photon impurity ansatz in Eq. (38) (purple dashed line) and the Gibbs state in Eq. (45) (black dotted line). Panels (b-e) shows the comparison between two cuts of $W(\alpha,\alpha^{*})$ [the one at $\textrm{Re}(\alpha)=0$ (blue curve) and the one at the one at $\textrm{Im}(\alpha)=0$ (orange curve)] and the fitted Wigner function (black dashed line). The values of $\ell$ are (b) $\ell=1$, (c) $\ell=200$, (d) $\ell=400$.
@@
In Fig. 11, we show the results of the fitting procedure corresponding to Fig. 5. In Fig. 11 (a), we plot the $L^{2}$ norm as a function of the index site $\ell$ for the chain with length $L=400$, using both the 2-photon decay and the generalized Scully-Lamb model as single-site ansätze. The results are compared with the $L^{2}$ norm corresponding to the fit of the thermal ansatz discussed in the following section. When $\ell\leq 50$ the fitting procedure does not give reliable results. In the bulk and in the right tail of the chain, instead, the local Wigner function is completely captured by both the single-site impurity models. In Figs. 11 (b-d) we show the comparison between $W_{\ell}(\alpha,\alpha^{*})$ and $W^{\rm fit}_{\ell}(\alpha,\alpha^{*})$. For $\ell=1$ [c.f. Fig. 11 (b)], the local Wigner function $W_{\ell}(\alpha,\alpha^{*})$ is not captured by the single-site ansatz. We conclude that within the left tail of the chain, in proximity of the coherent drive, different cavities remain correlated and thus a single-site description is not sufficient. For $\ell=200,400$ [c.f. Fig. 11 (c) and (d)] we see how $W_{\ell}(\alpha,\alpha^{*})$ matches with $W^{\rm fit}_{\ell}(\alpha,\alpha^{*})$.
@@
Here, we show that, in the prethermal phase of the chaotic Bose-Hubbard chain, the dephasing rate $\gamma^{\phi}$ is much larger than the 2-photon decay rate $\gamma^{\rm s}$ or the saturation $\mathcal{S}$, hence establishing the absence of long-lived phase coherence typical of lasing states. As discussed above, $\gamma^{\phi}$ cannot be extracted from the local steady-state Wigner function $W^{\rm ss}_{\ell}(\alpha,\alpha^{*})$. Here, we extract a rough estimate of $\gamma^{\phi}$ from two-time correlation functions in the prethermal domain in an $L=200$ chaotic chain by fitting them to a simple ansatz.
@@
Here, $A_{\ell,j}$ are real-valued amplitudes, $\Gamma_{\ell,j}\geq 0$ are decaying rates with magnitudes on the order of $\gamma^{\phi}$ [65], while the frequencies $\Omega_{\ell,j}$ capture possible oscillations in the early times. This simple ansatz successfully captures the auto-correlation function in the prethermal phase, as shown in Fig. 12 (black dashed lines). In the following table, we compare the resulting decay rates $\Gamma_{1}$ and $\Gamma_{2}$ with the saturation parameters obtained from the fit of $W^{\rm ss}_{\ell}(\alpha,\alpha^{*})$.
@@
In the three cases, the exponential decay rate of $C_{\ell}(\tau)$ is much larger than $\gamma^{\rm s}$ and $\mathcal{S}$, indicating the presence of a large dephasing rate $\gamma^{\phi}$, that is included in the 2-photon decay and generalized Scully-Lamb models. This analysis shows that the prethermal domain is not a laser since the latter would exhibit a significant phase coherence in time. Note that this distinction with a laser relies on the analysis of dynamical properties of the prethermal domain that go beyond the characterization of its static properties. This motivates our choice to describe the local physics of the chain in terms of a driven-dissipative impurity model rather than a Gibbs state ansatz (see the following section).
@@
Figure 13: Comparison between two cuts of $W(\alpha,\alpha^{*})$ [the one at $\textrm{Re}(\alpha)=0$ (blue curve) and the one at the one at $\textrm{Im}(\alpha)=0$ (orange curve)] and the fitted Wigner functions with the Gibbs state in Eq. (45), by imposing $U=0$ (green dashed line) and $U=0.1$ (black dash dotted line). Data are plotted in log scale on the y axis for a better visualization of the differences between the two fits. The drive strength is fixed to $F=7.5$ and the other parameters are set as in Fig. 3.
@@
As explained in the main text, the local statics can also be captured by a Gibbs state reading
$\hat{\rho}^{\rm eq}_{\ell}=\frac{\exp[-\beta_{\ell}^{\rm eff}(\hat{H}_{\rm imp}-\mu_{\ell}\,\hat{a}^{\dagger}a)]}{\operatorname{Tr}\left(\exp[-\beta_{\ell}^{\rm eff}(\hat{H}_{\rm imp}-\mu_{\ell}\,\hat{a}^{\dagger}a)]\right)}\,,$ (45)
@@
where the impurity Hamiltonian $\hat{H}_{\rm imp}$ is given by Eq. (8), $\beta^{\rm eff}_{\ell}$ is the effective temperature and $\mu_{\ell}$ is the chemical potential at the $\ell$-th site. By reabsorbing all the quadratic contributions into the effective chemical potential $\mu_{\ell}^{\rm eff}:=\mu_{\ell}-(\omega_{0}-U/2)$, we find that the local Wigner functions can simply be described by the two parameters $\beta_{\ell}^{\rm eff}$ and $\mu_{\ell}^{\rm eff}$. Note that, contrary to the driven-dissipative impurity models, the precise content of $\hat{H}_{\rm imp}$ is relevant for the fitting procedure to the Gibbs state. In particular, we fix $U=0.1$ [the value we set in all the numerical simulations of Eq. (2) presented in this work]. The $L^{2}$ norm associated to this fitting procedure is plotted in Fig. 11 (a). The result shows that the Gibbs state in Eq. (45) is successful in describing the local physics of the chain in the prethermal and thermal domains.
In Fig. 5 (a), we presented the product $\beta_{\ell}^{\rm eff}\mu_{\ell}^{\rm eff}$ for the three models considered in the paper, namely the 2-photon decay model in Eq. (38), the Scully-Lamb model in Eq. (39), and the Gibbs state in Eq. (45). All the proposed single-site ansätze give qualitatively similar results, thus validating the robustness of our local approach in capturing the features of $W_{\ell}^{\rm ss}(\alpha,\alpha^{*})$. A more sophisticated study of space and time correlation functions, and their characterization with multiple-site impurities, should be addressed in future studies.
@@
Finally, we show that taking into account the Kerr nonlinearity $U$ is important to the characterization of the local state in the thermal phase. In Fig. 13, we plot the Wigner function for $\ell=400$ at $\textrm{Im}(\alpha)=0$ and $\textrm{Re}(\alpha)=0$. We first perform the fit to the Gibbs state by imposing $U=0$ (black dashed line), which coincides with a Gaussian fit with a single parameter, $\beta_{\ell}^{\rm eff}\mu_{\ell}^{\rm eff}$. In this case, the fit does not reproduce the local Wigner functions. Instead, when we impose $U=0.1$ (gray dash-dotted line), the two-parameters ansatz perfectly reproduces the shape of $W_{\ell}^{\rm ss}(\alpha,\alpha^{*})$.

### v3

@@ 1. Numerical approaches
$\hat{L}_{\uparrow}=\hat{a}^{\dagger}(\gamma^{\uparrow}-\mathcal{S}\hat{a}\hat{a}^{\dagger})/\sqrt{\gamma^{\uparrow}},\qquad\hat{L}_{\phi}=\sqrt{\gamma^{\phi}}\hat{a}\hat{a}^{\dagger}=\sqrt{\eta+\frac{3\mathcal{S}}{2}}\hat{a}\hat{a}^{\dagger},\qquad\hat{L}_{\downarrow}=\sqrt{\gamma^{\downarrow}}\hat{a}.$ (42)
@@
while for the first model one requires $\gamma^{\rm s}\langle\hat{a}\hat{a}^{\dagger}\rangle\ll\gamma^{\uparrow}$. The choice of $\hat{H}_{\rm imp}$ and $\gamma^{\phi}$ does not affect the steady-state properties, including the steady-state Wigner function $W(\alpha,\alpha^{*})$. In the steady state, they have an effect only on time-correlation functions. In particular, by imposing $\gamma^{\phi}=3\mathcal{S}/2$ one finds oscillating time-correlation functions with a usually very small decay rate (that is $23\mathcal{S}/$ 2 which is by assumption much smaller than $2\gamma^{\uparrow}/\langle\hat{a}\hat{a}^{\dagger}\rangle$). If instead $\gamma^{\phi}\gg\mathcal{S}$ one obtains the generalized version of the Scully-Lamb model exhibiting rapidly decaying time-correlation functions in the lasing phase. A similar phenomenology occurs in the 2-photon decay impurity model, where the phase coherence is suppressed if $\gamma^{\phi}\gg\gamma^{\rm s}$.
While the two impurity models above exhibits very similar steady-state properties, a finite effective temperature $T$ can, in principle, only be straightfowardly associated with the 2-photon decay model. Indeed, the set of Lindblad operators in Eq. (41) can be easily interpreted as coupling to a set of thermal reservoirs. $\hat{L}_{\rm s}$ corresponds to a zero-temperature 2-photon reservoir. $\hat{L}_{\uparrow}$ and $\hat{L}_{\downarrow}$ correspond to finite-temperature 1-photon reservoir at a temperature set by the detailed balance
$\frac{\gamma^{\uparrow}}{\gamma^{\downarrow}}=:{\rm e}^{-(\omega_{0}-\mu)/T}=:{\rm e}^{\mu/T},\quad\textit{i.e.}\quad\mu/T:=\log\left({\gamma_{\uparrow}}\big{/}{\gamma_{\downarrow}}\right),$ (44)
@@
Figure 15: Fit of the $\ell$-th site Wigner function of the driven-dissipative bosonic chain described by Eqs. (2) and (1) for a chain with length $L=400$. In panel $(a)$, we show the $L^{2}$ norm in Eq. (45) between the $L$-th site Wigner function and the fitted Wigner function obtained from the three ansätze considered in the article: the Gibbs state in Eqs. (8) and (9) (green line), the Scully-Lamb model in Eq. (42) (purple-dashed line), and the 2-photon driven-dissipative impurity model in Eq. (41) (red-dashed line). Panels (b-e) shows the comparison between two cuts of $W(\alpha,\alpha^{*})$ [the one at $\textrm{Re}(\alpha)=0$ (blue curve) and the one at the one at $\textrm{Im}(\alpha)=0$ (orange curve)] and the fitted Wigner function (black dashed line). The values of $\ell$ are (b) $\ell=1$, (c) $\ell=200$, (d) $\ell=400$.
@@
In Fig. 15, we show the results of the fitting procedure corresponding to Fig. 4. In Fig. 15 (a), we plot the $L^{2}$ norm as a function of the index site $\ell$ for the chain with length $L=400$, using both the 2-photon decay and the generalized Scully-Lamb model as single-site ansätze. The results are compared with the $L^{2}$ norm corresponding to the fit of the thermal ansatz discussed in the following section. When $\ell\leq 50$ the fitting procedure does not give reliable results. In the bulk and in the right tail of the chain, instead, the local Wigner function is completely captured by both the single-site impurity models. In Figs. 15 (b-d) we show the comparison between $W_{\ell}(\alpha,\alpha^{*})$ and $W^{\rm fit}_{\ell}(\alpha,\alpha^{*})$. For $\ell=1$ [c.f. Fig. 15(b)], the local Wigner function $W_{\ell}(\alpha,\alpha^{*})$ is not captured by the single-site ansatz. We conclude that within the left tail of the chain, in proximity of the coherent drive, different cavities remain correlated and thus a single-site description is not sufficient. For $\ell=200,400$ [see Fig. 15 (c) and (d)] we see how $W_{\ell}(\alpha,\alpha^{*})$ matches with $W^{\rm fit}_{\ell}(\alpha,\alpha^{*})$.
@@
Here, we show that, in the prethermal phase of the chaotic Bose-Hubbard chain, the dephasing rate $\gamma^{\phi}$ is much larger than the 2-photon decay rate $\gamma^{\rm s}$ or the saturation $\mathcal{S}$, hence establishing the absence of long-lived phase coherence typical of lasing states. As discussed above, $\gamma^{\phi}$ cannot be extracted from the local steady-state Wigner function $W_{\ell}(\alpha,\alpha^{*})$. Here, we extract a rough estimate of $\gamma^{\phi}$ from two-time correlation functions in the prethermal domain in an $L=200$ chaotic chain by fitting them to a simple ansatz.
@@
Here, $A_{\ell,j}$ are real-valued amplitudes, $\Gamma_{\ell,j}\geq 0$ are decaying rates with magnitudes on the order of $\gamma^{\phi}$ [83], while the frequencies $\Omega_{\ell,j}$ capture possible oscillations in the early times. This simple ansatz successfully captures the auto-correlation function in the prethermal phase, as shown in Fig. 16 (black dashed lines). In the following table, we compare the resulting decay rates $\Gamma_{1}$ and $\Gamma_{2}$ with the saturation parameters obtained from the fit of $W_{\ell}(\alpha,\alpha^{*})$.
@@
In the three cases, the exponential decay rate of $C_{\ell}(\tau)$ is much larger than $\gamma^{\rm s}$ and $\mathcal{S}$, indicating the presence of a large dephasing rate $\gamma^{\phi}$, that is included in the 2-photon decay and generalized Scully-Lamb models. This analysis shows that the prethermal domain is not a laser since the latter would exhibit a significant phase coherence in time. Note that this distinction with a laser relies on the analysis of dynamical properties of the prethermal domain that go beyond the characterization of its static properties. This motivates our choice to describe the local physics of the chain in terms of a driven-dissipative impurity model and not only the Gibbs state ansatz (see the following section).
@@
As explained in the main text, in the chaotic regime the local statics can be captured by a Gibbs state reading
$\hat{\rho}^{\rm eq}_{\ell}=\frac{\exp[-(\hat{h}-\mu_{\ell}\,\hat{a}^{\dagger}\hat{a})/T_{\ell}]}{\operatorname{Tr}\left(\exp[-(\hat{h}-\mu_{\ell}\,\hat{a}^{\dagger}\hat{a})/T_{\ell}]\right)}\,,$ (48)
@@
where the impurity Hamiltonian $\hat{h}$ is given by Eq. (9), $T_{\ell}$ is the effective temperature and $\mu_{\ell}$ is the chemical potential at the $\ell$-th site. By reabsorbing all the onsite quadratic contributions into the effective chemical potential $\mu_{\ell}\leftarrow\mu_{\ell}-(\omega_{0}-U/2)$, we find that the local Wigner functions can simply be described by the two parameters $T_{\ell}$ and $\mu_{\ell}$. Note that, contrary to the driven-dissipative impurity models, the precise content of $\hat{h}$ is relevant for the fitting procedure to the Gibbs state. In particular, we fix $U=0.1$ [the value we set in all the numerical simulations of Eq. (2) presented in this work]. The $L^{2}$ norm associated to this fitting procedure is plotted in Fig. 15 (a). The result shows that the Gibbs state in Eq. (48) is successful in describing the local physics of the chain in the prethermal and thermal domains. In Fig. 4 (d), we presented the dimensionless quantity $\mu_{\ell}/T_{\ell}$ for the three models considered in the paper, namely the 2-photon decay model in Eq. (41), the Scully-Lamb model in Eq. (42), and the Gibbs state in Eq. (48). All the proposed single-site ansätze give qualitatively similar results, thus validating the robustness of our local approach in capturing the features of $W_{\ell}(\alpha,\alpha^{*})$. A more sophisticated study of space and time correlation functions, and their characterization with multiple-site impurities, should be addressed in future studies.
@@
Finally, we show that taking into account the Kerr nonlinearity $U$ is important to the characterization of the local state in the chaotic phase at intermediate drive amplitude, even in the thermal domain, while it is not relevant in the chaotic phase at low drive amplitude, where the thermal domain extends across the whole chain and the equipartition theorem describes the local distribution of momenta. In Fig. 17 (a), we plot the Wigner function for $\ell=370$ at $\textrm{Im}(\alpha)=0$ (blue curve) and $\textrm{Re}(\alpha)=0$ (orange curve) and $F=5.5$. We perform the fit to the Gibbs state by imposing $U=0$ (black dashed line), which coincides with a Gaussian fit with a single parameter, $\mu_{\ell}/T_{\ell}$. The Gaussian ansatz reproduces the shape of the Wigner function, thus further validating the use of the equipartition theorem in this regime. In Fig. 17 (b), we plot the Wigner function for $\ell=370$ at $\textrm{Im}(\alpha)=0$ (blue curve) and $\textrm{Re}(\alpha)=0$ (orange curve) and $F=7.5$. We first perform the the single-parameter Gaussian fit. In this case, the Gaussian Ansatz does not reproduce the local Wigner functions. Instead, when we impose $U=0.1$ (gray dash-dotted line), the full Gibbs state perfectly reproduces the shape of $W_{\ell}(\alpha,\alpha^{*})$. This proves that the contribution of the nonlinearity at intermediate drive amplitudes is present not only in the prethermal domain, where the nonlinearity determines the saturation of the phton number, but also in the thermal domain, where the vacuum dressed by thermal fluctuations accounts for a small but finite amount of nonlinearity.
Figure 17: Comparison between two cuts of $W(\alpha,\alpha^{*})$ and the fitted Wigner functions with the Gibbs state in Eq. (48). (a) Comparison between $W_{\ell}[\textrm{Re}(\alpha=0)]$ (blue curve) and $W_{\ell}[\textrm{Im}(\alpha=0)]$ (orange curve) and the fitted Wigner function for $\ell=370$ in a $L=400$ chain at low drive amplitude $F=5.5$, by imposing $U=0$ (black-dashed line). (b) Comparison between $W_{\ell}[\textrm{Re}(\alpha=0)]$ (blue curve) and $W_{\ell}[\textrm{Im}(\alpha=0)]$ (orange curve) and the fitted Wigner function for $\ell=370$ in a $L=400$ chain at intermediate drive amplitude $F=7.5$, by imposing $U=0.1$ (black-dashed line) and $U=0$ (red line) and $U=0.1$ (black dash dotted line). The other parameters are set as in Fig. 2.
Additional details on the RNW regime
In Sec. II.4 we studied extensively the RNW regime, highlighting its distinctive quantum features, encoded in the phase fluctuations at the same site $\ell$, measured by means of the circular variance $\Delta\varphi_{\ell}:=1-|\langle{\rm e}^{{\rm i}\varphi_{\ell}}\rangle|$, and at different sites $k$ and $\ell$, measured by means of the first-order coherence function $g^{(1)}_{k\ell}$ in Eq. (7). We present here three additional observables, the photon number $n_{\ell}$, the average phase $\langle\varphi_{\ell}\rangle$, and the bosonic field $\alpha_{\ell}$. The comparison between the classical and TWA solutions is presented in Fig. 18.
We find that both the photon number and the averaged phase coincide when estimated with the TWA and the classical approach, as shown in Fig. 18 (a) and (b). For the bosonic field $\alpha_{\ell}$, reported in Fig. 18 (c), we clearly observe the effect of quantum fluctuations, which causes the decay of $\textrm{Re}(\alpha_{\ell})$ with $\ell$ [the same pattern can be observed for $\textrm{Im}(\alpha_{\ell})$ and $|\alpha_{\ell}|$]. This distinction can be explained as follows: quantum fluctuations induce an angular spreading of the Wigner function [see Figs. 6 (c-f) in the main text] but keeping a radial localization around the classical value. As a result, the average $\langle|\alpha_{\ell}|^{2}\rangle$ only encodes the information about the radius of $W_{\ell}(\alpha,\alpha^{*})$, which coincides with the classical value, while the phase spreads symmetrically around the classical value, and its average again returns the Gross-Pitaevskii solution. An increasing $\Delta\varphi_{\ell}$ leads to the decay of $\alpha_{\ell}$ because the latter quantity encodes both the radius and the angular spreading of $W_{\ell}(\alpha,\alpha^{*})$, resulting in a decreasing $\textrm{Re}(\alpha_{\ell})$ when the average is performed.
Figure 18: Additional comparison between the classical and TWA solutions for the RNW regime. (a) Photon number $n_{\ell}$ as a function of $\ell$ for a $L=100$ chain in the RNW regime. The blue curve is are the TWA values whereas the black-dashed curve represents the classical Gross-Pitaevskii prediction. (b) Same as in (a) but for the averaged phase $\langle\varphi_{\ell}\rangle$. (c) Same as in (a) but for the real part of the bosonic field, $\textrm{Re}(\alpha_{\ell})$. Results are computed by averaging over $N_{\rm traj}=5\times 10^{3}$ independent Wigner trajectories and over a time window $\Delta\tau=10^{3}$ once the steady state is reached. The drive amplitude is set to $F=12.5$. The other parameters are set as in Fig. 2.
Effects of an intrinsic dissipation rate
Figure 19: Effects of an instrinsic dissipation rate $\gamma_{\rm int}$ on the chaotic and the RNW regime. (a) photon number $n_{\ell}$ as a function of $\ell$ for a $L=200$ chain in the chaotic regime at $F=7.5$ for various intrinsic dissipation rates $\gamma_{\rm int}=10^{-4},10^{-3},10^{-2},10^{-1}$ (from dark purple to light pink). (b-e) Local Wigner function $W_{\ell}\left(\alpha,\alpha^{*}\right)$ in the chaotic chain at $\ell=50$ for the same values of $\gamma_{\rm int}$ considered in panel (a). (f) Same as in panel (a) but for a $L=100$ chain in the RNW regime at $F=12.5$. (g-j) Same but for the regular RNW steady state at $\ell=50$. The other parameters are set as in Fig. 2.
The setup described by Eqs. (1) and (2) presents a notable idealization: the intrinsic loss of each resonator is assumed to be zero. In arrays of superconducting nonlinear resonators [43, 44], a finite intrinsic dissipation rate $\gamma_{\rm int}\ll\gamma$ is always present. Therefore, Eq. (1) generalizes to
$\frac{\partial\hat{\rho}}{\partial t}$ $=-{\rm i}[\hat{H},\hat{\rho}]+\mathcal{D}[\hat{L}_{1}]\hat{\rho}+\mathcal{D}[\hat{L}_{L}]\hat{\rho}+\sum_{\ell=1}^{L}\mathcal{D}[\hat{L}_{\ell}^{\rm int}]\hat{\rho},$ (49)
where $\hat{L}^{\rm int}_{\ell}=\sqrt{\gamma_{\rm int}}\hat{a}_{\ell}$. An important question concerns the robustness of the classical and quantum photonic regimes studied in the main text with respect to spurious intrinsic dissipations. This study is relevant for the experimental realization of the driven-dissipative architecture and the observation of the phenomena we describe. Here, we focus on the effects of a finite $\gamma_{\rm int}$ on the chaotic prethermal domain in a $L=200$ chain at $F=7.5$, and on the regular RNW regime in a $L=100$ chain at $F=12.5$.
In Fig. 19 (a) we plot the photon number $n_{\ell}$ as a function of the site index $\ell$ in the chaotic phase. For small dissipation rates equal to $\gamma_{\rm int}=10^{-4}$ the photon number profile resembles the one presented in the main text [see Fig. 3 (a)]. For larger dissipation rates equal to $\gamma_{\rm int}=10^{-3}$ the prethermal phase visibly shrinks. For even larger dissipation rates up to $\gamma_{\rm int}=10^{-1}$, the prethermal regime disappears completely and most of the resonator end in a vacuum-like state with around zero photons. This discussion is supported by the plots of the local Wigner functions $W_{\ell}\left(\alpha,\alpha^{*}\right)$ reported in Figs. 19 (b-e) for $\ell=50$. Similar considerations hold for the regular RNW regime. In Fig. 19 (f) we plot the photon number $n_{\ell}$ as a function of the site index $\ell$. For small dissipation rates equal to $\gamma_{\rm int}=10^{-4}-10^{-3}$ we observe flat photon number profile of the RNW regime [see Fig. 5 (a) and Fig. 18 (a)]. For larger intrinsic dissipation rates, like $\gamma_{\rm int}=10^{-2}$, the photon number acquires a gradient with $\ell$ and the RNW regime is destroyed. For even larger intrinsic dissipation rates up to $\gamma_{\rm int}=10^{-1}$, the bulk of the chain goes into a vacuum-like states with few or zero photons in each resonator. This discussion is supported by the plots of the local Wigner functions $W_{\ell}\left(\alpha,\alpha^{*}\right)$ reported in Figs. 19 (g-j) for $\ell=50$.
We conclude that an intrinsic dissipation rate, reasonably, threatens the out-of-equilibrium regimes we studied in the main text. Too large values of $\gamma_{\rm int}$ prevent photon transport across the chain and most of the system is in a trivial vacuum-like state. On the other hand, both the prethermal and the RNW regimes are still present with a small but finite $\gamma_{\rm int}$.